OpenStudy (anonymous):

which equation is equivalent to 2(x-5)+7x+10(3x)=524 a)22x-5=524 b)22x-10=524 c)39x-5=524 d)39x-10=524

jimthompson5910 (jim_thompson5910):

2(x-5)+7x+10(3x)=524 2(x)+2(-5)+7x+10(3x)=524 2x-10+7x+30x=524 39x-10=524 So the answer is choice D